OpenStudy (anonymous):

(hºg)(-5) show work please. that is not a deg sign btw...not sure what it is but its a small "o" inbetween

OpenStudy (anonymous):

You need to know h(x) and g(x). The symbol is a composition symbol. It asks you to take h(g(x)).

OpenStudy (anonymous):

What are h(x) and g(x)?

OpenStudy (anonymous):

oh sorry @albertedison98 lol g(x)=2x h(x)=x^2+4

OpenStudy (anonymous):

So you want to know h(g(-5)). Start with the inside: g(-5)=2(-5)=-10. Then, h(-10)=(-10)^2+4=100+4=104. So, h(g(-5)=104.

OpenStudy (anonymous):

thanks so much @albertedison98 !!!!!!